The base of a ladder is 14 feet away from the wall. The top of the ladder is 17 feet from the floor. Find the length of the ladder to the nearest thousandth.

Answer by checkley75(3666)

A^2+B^2=C^2
14^2+17^2=C^2
196+289=C^2
485=C^2
C=SQRT485
C=22.02 FEET. ANSWER FOR THE LENGTH OF THE LADDER.